OpenStudy (anonymous):

How do I solve the quadratic equation by factoring and setting each factor equal to zero in this problem : 7d(squared) + 18d +8 = 0

OpenStudy (the_fizicx99):

Hm, 7d^2 + 18d + 8 = 0 7 * 8 = 56 14 * 4 = 56 14 + 4 = 18 7d^2 + 14d + 4d + 8 = 0 Factor by grouping, take the first two terms 7d^2 + 14d, they have a GCF in both the coefficient and variable. 7 and d 7d(d + 2) 4d + 8 only has 4, 4(d + 2) (7d + 4)(d + 2) = 0 solve for D, 7d + 4 = 0 -4 -4 7d = -4 /7 /7 D = -4/7 is one solutions D + 2 = 0 -2 -2 D = -2 is your other solution. When you substitute in -2 or -4/7 you'll get 0